Average Speed 3 Puntos:1 Due Vie, 25. En 2019, 22:30:00 hs (AST) 0 1 23 4 5 67 8 9 10 11 12 13 In this graph of displacement versus time, what is the average speed in going from point F to point J in m/s? (1/4) Hint: Find the distance traveled in each section and add them, then divide by the time elapsed Enviar Respuesta Sólo se requiere un número. Computadora lee la unidades de /4). Tries 5/10 Intentos Anteriores Enviar Mensaje de Discusión En
Answer #1

distance covered in FG = 3-2 =1

distance covered in GH = 3-3 =0

distance covered in HI = |1-3| = 2
distance covered in IJ = 0

total distance travelled = 3

average speed = 3/4 = 0.75 m/s

please rate it up thanks :)